OpenStudy (anonymous):

Id F(x)=3x-9 and g(x) = -2x+2, find (f(times)g) (x)

OpenStudy (anonymous):

*if

OpenStudy (hoblos):

(3x-9)(-2x+2)= -6x^2 +24x-18

OpenStudy (anonymous):

=3(-2x+2)-9 =-6x+6-9 =-6x-3 it is a composite function

OpenStudy (anonymous):

They are both answer choices

OpenStudy (anonymous):

first one is wrong